### v1

# Hypertangency of plane curves and the algebraic exceptional set

Abstract.
We investigate plane curves intersecting in at most two unibranched points to study the algebraic exceptional set appearing in standard conjectures of diophantine and hyperbolic geometry. Our first result compares the local geometry of two hypertangent curves, i.e. curves having maximal contact at one unibranched point. This is applied to fully describe the exceptional set and, more generally, the hyper-bitangency set, of a plane curve with three components.
Key words and phrases:
Exceptional set, algebraic curves, hypertangency, Vojta conjecture.
2010 Mathematics Subject Classification:
14H20, 14H45, 14G40
1. Introduction
The goal of this paper is to explore some geometric properties of plane curves that intersect a given curve in few points.
The main motivation comes from a series of conjectures of Lang, Vojta Demailly and Campana, as in [Lan91, Conjecture 3.7], [Voj87, Conjecture 3.4.3] [Dem97] and [Cam04, Conjectures 9.2 and 9.20] (see [BG06, HS00, AT20, DT15] for introductions and discussions), on the distribution of integral and rational points on quasi-projective surfaces defined over number fields and function fields. For our goals we state here a consequence of their more general conjectures.
Conjecture 1.
Let $B\subset\mathbb{P}^{2}$ be a reduced curve with at most simple normal crossing singularities (i.e. nodes). If $\deg B\geq 4$, then the set of rational curves $C\subset\mathbb{P}^{2}$ such that $|\nu_{C}^{-1}(C\cap B)|\leq 2$ is finite, where $\nu_{C}:C^{\nu}\to C$ is the normalization.
The above set is often referred to as the algebraic (or geometric) exceptional set in the aforementioned circle of conjectures (see for example [Lan86]); we denote it by $\mathcal{E}(B)$, i.e.
(1) $\mathcal{E}(B):=\{C\subset\mathbb{P}^{2}:\ C\text{ rational},\ \ |\nu_{C}^{-1}(C\cap B)|\leq 2\}.$
Conjecturally, $\mathcal{E}(B)$ is made of the curves that contain the only potentially infinite families of integral points in the affine surface $\mathbb{P}^{2}\setminus B$. We stress here that the condition “degree of $B$ at least 4” is equivalent to the fact that $(\mathbb{P}^{2},B)$ is of log general type.
Historically, describing the distribution of integral points (and the exceptional set) in an affine surface of log-general type of the form $\mathbb{P}^{2}\setminus B$, as predicted by the conjectures of Lang and Vojta, has proven harder the fewer the irreducible components of the curve $B$. For instance, if $B$ has at least 4 components, much is known of the surface $\mathbb{P}^{2}\setminus B$, and in particular the above mentioned conjectures follow as an application of Schmidt’s Subspace Theorem. The analogue result over function fields is a consequence of a generalized version of Mason-Stother’s abc theorem for polynomials, see [BM86].
When $B$ has three irreducible components, Vojta’s Conjecture over number field is still an open problem. On the other hand, the proof of the function field version of Vojta’s Conjecture in [CZ08] and [CZ13], for the surface $\mathbb{P}^{2}\setminus B$ when $B$ has at least three irreducible components, implies that in this case all the curves in $\mathcal{E}(B)$ have bounded degree (the same result in the so-called non-split case was proved by the first author together with Capuano in [CT22] building on the previous work [Tur17]). More recently [GNSW23] showed (as a special case of their Theorem 4) that, in the same settings, the exceptional set $\mathcal{E}(B)$ is indeed a closed subset of $\mathbb{P}^{2}$ and described a closed set containing it. Similar results had been previously obtained by Corvaja and Zannier in [CZ00].
In another direction one can look at the very general situation, i.e. when the divisor $B$ is very general in the space of all curves of given degree. In this setting, Chen, Riedl and Yeong in [CRY23] showed that, when $B$ has degree 4 and it is very general (in particular smooth irreducible), the exceptional set $\mathcal{E}(B)$ consists only of the bitangent lines and the flex lines, answering a question of Lang. On the other hand, when $B$ has degree at least 5 and it is very general, the exceptional set $\mathcal{E}(B)$ is empty: this has been proven independently in [Che04] and in [PR07]. No similar result is known without the very general assumption.
In the present article we focus on the case where $B$ has (at least) three irreducible components. We describe the exceptional set as a consequence of Theorem 3.3.2 where we we show that for a general $B$, an integral curve $C$ of arbitrary genus, such that $|\nu_{C}^{-1}(C\cap B)|\leq 2$, has degree at most 2 (where again $\nu_{C}$ is the normalization). Moreover we give sharp upper bounds for the number of conics and lines and characterize precisely the non general cases. To our knowledge the independence from the genus is a new feature that was not observed before.
Following Lang’s philosophy, Conjecture 1 is related to the hyperbolicity of the surface $\mathbb{P}^{2}\setminus B$. In this setting, the Logarithmic Kobayashi Conjecture predicts that for a general $B$ of degree at least five, the surface $\mathbb{P}^{2}\setminus B$ is Brody hyperbolic, i.e. every holomorphic map $\mathbb{C}\to\mathbb{P}^{2}\setminus B$ is constant. This implies in particular that there are no copies of $\mathbb{C}$ or $\mathbb{C}^{*}$ in the complement of $B$. Therefore, there cannot exist any rational curves $C$ that meet $B$ in at most two unibranched points (for in this case one would get the non constant map $\exp:\mathbb{C}\to\mathbb{C}^{*}\subset\mathbb{P}^{2}(\mathbb{C})\setminus B$). In a parallel direction, the logarithmic Green-Griffiths-Lang Conjecture predicts that, when $\deg B\geq 4$, there exists an analytic exceptional set, i.e. a proper closed subset of $\mathbb{P}^{2}\setminus B$ that contains all the images of the non constant holomorphic maps $\mathbb{C}\to\mathbb{P}^{2}\setminus B$. Conjecturally the analytic exceptional set and the algebraic exceptional set should coincide, as predicted by Lang; nevertheless the analytic exceptional set always contains the algebraic one. Therefore our main results can be seen as partial progress in the description of the analytic exceptional set for $\mathbb{P}^{2}\setminus B$, in the case where $B$ has at least three irreducible components. In this latter case, is worth mentioning that the corresponding analytic result of [CZ08] and [CZ13] was obtained in [NWY08]
To state our results more precisely we introduce the set of curves “hyper-bitangent” to $B$
$\operatorname{Hyp}(B,2):=\{C\subset\mathbb{P}^{2}:\ C\text{ integral},\ \ |\nu_{C}^{-1}(C\cap B)|\leq 2\}$
so that, of course, $\mathcal{E}(B)\subset\operatorname{Hyp}(B,2)$. We summarize in one statement our main results; see Proposition 3.2.1, and Theorems 3.3.1, 3.3.2, and 3.3.3 for more precise and stronger statements.
Theorem.
Let $B\subset\mathbb{P}^{2}$ be a reduced curve with three irreducible components and only nodal singularities. If $\deg B\geq 4$ then
  1. (a)
    $\mathcal{E}(B)=\operatorname{Hyp}(B,2)$;
  2. (b)
    $\operatorname{Hyp}(B,2)$ is finite (and effectively bounded);
  3. (c)
    if $\deg B\geq 5$ and $B$ is general, then $\operatorname{Hyp}(B,2)$ is empty.
The proof of these facts uses a geometric result of independent interest, Theorem 2.2.1, which analyzes the local geometry of two plane curves, $B$ and $C$, meeting in one point, i.e. $C\cap B=\{q\}$. The precise statement of this theorem requires some technical preliminaries; informally speaking, it establishes that the local geometries of $B$ and $C$ at the point $q$ are closely related.
Outline of the paper.
In Section 2 we establish our principal geometric tool, Theorem 2.2.1, on the local geometry of two plane curves meeting in only one unibranched point. In Section 3 we prove our main results describing the sets $\operatorname{Hyp}(B,2)$ for curves with three components, thus proving the previously stated Theorem. Finally, Section 4 collects various special cases and examples related to the earlier topics.
Notation
We work over $\mathbb{C}$. We denote by $C$ an integral (i.e. reduced and irreducible) projective curve lying in a smooth projective rational surface $S$. We denote by $\nu_{C}:C^{\nu}\to C$ the normalization, by $p_{a}(C)$ the arithmetic genus and by $g(C)=p_{a}(C^{\nu})$ the geometric genus of $C$. A rational curve is an integral curve of geometric genus zero. Given a point $p\in C$, we write ${\operatorname{mult}}_{p}(C)$ for the multiplicity of $C$ at $p$. We say that $p$ is unibranched if $|\nu_{C}^{-1}(p)|=1$. We denote by $C^{\nu}_{p}$ the partial normalization of $C$ at $p$; the so-called $\delta$-invariant, $\delta_{C}(p)$, of $p$ is $\delta_{C}(p)=p_{a}(C)-p_{a}(C^{\nu}_{p}).$ The following formula is well-known
(2) $\delta_{C}(p)=\sum m_{q}(m_{q}-1)/2$
where $q$ varies over all points infinitely near to $p$ (including $p$), and $m_{q}$ is the multiplicity of $q$.
If $C,B\subset S$ are reduced curves with no components in common, and $p\in C\cap B$, we write $(C\cdot B)_{p}$ for their multiplicity of intersection at $p$. We say that $C$ is hypertangent to $B$ if $|\nu_{C}^{-1}(B\cap C)|=1$, so that also $|B\cap C|=1$.
Let now $S=\mathbb{P}^{2}$, fix a reduced curve $B\subset\mathbb{P}^{2}$ and a point $q\in B$. The set of integral curves hypertangent to $B$ at $q$ is denoted as follows
$\operatorname{Hyp}(B;q)=\{C\subset\mathbb{P}^{2}:\ C\text{ integral},\ C\cap B=\{q\},\ |\nu_{C}^{-1}(q)|=1\},$
and the set of all integral curves hypertangent to $B$ is
$\operatorname{Hyp}(B,1):=\cup_{q\in B}\operatorname{Hyp}(B;q).$
For a positive integer $d$ we write $\operatorname{Hyp}_{d}(B;q)\subset\operatorname{Hyp}(B;q)$ and $\operatorname{Hyp}_{d}(B,1)\subset\operatorname{Hyp}(B,1)$, for the subsets parametrizing curves of degree $d$; we view both of them as subspaces of the projective space $\mathbb{P}^{d(d+3)/2}$ parametrizing plane curves of degree $d$.
If $C\in\operatorname{Hyp}_{d}(B;q)$ and $C$ is singular at $q$, then it necessarily has a unibranched singularity. We set for $m\geq 1$
$\operatorname{Hyp}^{m}(B;q)=\{C\in\operatorname{Hyp}(B;q):\ {\operatorname{mult}}_{q}(C)=m\}$
and we define $\operatorname{Hyp}_{d}^{m}(B;q)\subset\operatorname{Hyp}_{d}(B;q)$ analogously.
Now we extend to double intersections. As we said before, $C$ is hyper-bitangent to $B$ if $\nu_{C}^{-1}(C\cap B)\leq 2$ and we denote by $\operatorname{Hyp}(B,2)$ the set of curves hyper-bitangent to $B$. We have $\operatorname{Hyp}(B,1)\subset\operatorname{Hyp}(B,2)$, i.e. hypertangents are special hyper-bitangents. We set
$\operatorname{Hyp}_{d}(B,2):=\{C\in\operatorname{Hyp}(B,2):\ \deg C=d\}$
so that $\operatorname{Hyp}_{d}(B,2)\subset\mathbb{P}^{d(d+3)/2}$.
Acknowledgements.
We are pleased to thank Laura Capuano, Ciro Ciliberto, Pietro Corvaja, Kristin DeVleming, Edoardo Sernesi and Umberto Zannier for discussions and useful comments about this work. LC is partially supported by PRIN 2017SSNZAW and PRIN 2022L34E7W, Moduli spaces and birational geometry. AT is partially supported by PRIN 2022HPSNCR: Semiabelian varieties, Galois representations and related Diophantine problems and PRIN 2020KKWT53: Curves, Ricci flat Varieties and their Interactions, and is a member of the INDAM group GNSAGA.
2. Hypertangency
2.1. Preliminaries on unibranched points
We here state some well known facts about unibranched points of curves; we refer to [Wal04] for an exhaustive treatise.
Let $q$ be a unibranched point of an integral curve $C\subset S$ lying on a smooth rational surface $S$. Our analysis here is local, therefore we are free to replace $S$ with an open neighborhood of $q$ isomorphic to $\mathbb{A}^{2}$. As $q$ is unibranched, the tangent line to $C$ at $q$, sometimes called the “tangent direction”, is locally well defined; we denote it by $L$. We say that $q$ is an $(m,n)$-point if ${\operatorname{mult}}_{q}(C)=m$ and if the multiplicity of the intersection of $L$ and $C$ in $q$ satisfies
$(C\cdot L)_{q}=n.$
If $m=1$ and $n\geq 3$ we say that $q$ is a flex of $C$.
We choose local coordinates, $x,y$, so that $q=(0,0)$, the line $L$ has equation $y=0$, and $C$ has equation $f(x,y)=0$ with $\deg f=d\geq n$ with
(3) $f(x,y)=a_{0,m}y^{m}+\sum_{m+1\leq i+j\leq d}a_{i,j}x^{i}y^{j}$
such that $a_{0,m}\neq 0$ and the smallest power of $x$ appearing in $f$ is $x^{n}$, i.e. $a_{i,0}=0$ for $i<n$ and $a_{0,n}\neq 0$. Notice that, as will be clear in the sequel, having local equation of type (3) is not sufficient for $q$ to be unibranched.
Consider the blow-up of $S$ at $q$, denote by $E$ the exceptional divisor and by $C^{\prime}$ and $L^{\prime}$ the strict transforms of $C$ and $L$. The map $\sigma:C^{\prime}\to C$ induced by this blow-up is bijective; set $q^{\prime}:=\sigma^{-1}(q)$. We state some known facts in a convenient form; we include the proof for completeness.
Lemma 2.1.1.
Let $0<m<n$ and let $q\in C\subset S$ be an $(m,n)$-point.
  1. (a)
    If $n<2m$ then $q^{\prime}$ is an $(n-m,m)$-point of $C^{\prime}$, and $C^{\prime}$ is tangent to $E$ at $q^{\prime}$.
  2. (b)
    If $n>2m$ then $q^{\prime}$ is an $(m,n-m)$-point of $C^{\prime}$, and $C^{\prime}$ is tangent to $L^{\prime}$ at $q^{\prime}$.
  3. (c)
    If $n=2m$ then $q^{\prime}$ is an $m$-fold point of $C^{\prime}$, and $C^{\prime}$ is neither tangent to $E$ nor to $L^{\prime}$ at $q^{\prime}$.
Proof.
To blow-up at $q$ we set $y=vx$ and use $(x,v)$ as local coordinates in the blow-up at $q^{\prime}=(0,0)$, the local equation of the exceptional divisor $E$ is $x=0$, and the local equation of $L^{\prime}$ is $v=0$. Let $f^{\prime}(x,v)=0$ be the affine equation of $C^{\prime}$ obtained from (3):
$f^{\prime}(x,v)=x^{-m}f(x,vx)=a_{0,m}v^{m}+\sum_{m+1\leq i+j\leq d}a_{i,j}x^{i+j-m}v^{j}.$
The smallest power of $x$ appearing as a summand of $f^{\prime}$ is $x^{n-m}$, and the smallest power of $v$ is $v^{m}$. We have three cases.
Case (a) $n<2m$. Hence $n-m<m$ and we set $r=n-m$. Since $C^{\prime}$ is unibranched at $q^{\prime}$, and $f^{\prime}$ contains the summand $x^{r}$ but not the summand $v^{r}$ (as $r<m$), all terms in $f^{\prime}$ of degree at most $r$ divisible by $xv$ must vanish (for otherwise the lowest homogeneous part of $f^{\prime}$ will be reducible, contradicting the fact that $q^{\prime}$ is a unibranched point of $C^{\prime}$). Hence the tangent line to $C^{\prime}$ at $q^{\prime}$ has local equation $x=0$, hence $C^{\prime}$ is tangent to $E$. Since the smallest power of $v$ in $f^{\prime}$ is $v^{m}$, we get that $q^{\prime}$ is an $(r,m)$-point. (a) is proved.
Case (b) $n>2m$. The smallest power of $v$ appearing in $f^{\prime}$ is $v^{m}$ and, arguing as in the previous case, $f^{\prime}$ has no other term of degree at most $m$. Hence the tangent line to $C^{\prime}$ at $q^{\prime}$ has local equation $v=0$, so that $C^{\prime}$ is tangent to $L$. As the smallest power of $x$ is $x^{n-m}$, we get that $q^{\prime}$ is an $(m,n-m)$-point. (b) is proved.
Case (c) $n=2m$. We have both $x^{n-m}=x^{m}$ and $v^{m}$ appearing in $f^{\prime}$ as smallest powers of $x$ and $v$. Since $C^{\prime}$ is unibranched at $q^{\prime}$, this is possible only if the homogeneous part,$f^{\prime}_{m}$, of degree $m$ of $f^{\prime}$, has form
$f^{\prime}_{m}=(\alpha v+\beta x)^{m}$
with $\alpha,\beta\neq 0$. Hence $q^{\prime}$ is an $m$-fold point, and the tangent line to $C^{\prime}$ there has local equation $\alpha v+\beta x=0$, hence it is different from $E$ and $L^{\prime}$. ∎
We will use the following well known facts.
Remark 2.1.2.
Let $B,C\subset S$ be two integral curves and $q\in B\cap C$. Suppose that $B$, respectively $C$, has a unibranched $m_{B}$-fold, resp. $m_{C}$-fold, point at $q$. If $\sigma:S^{\prime}\to S$ is the blow up at $q$ and $C^{\prime},B^{\prime}$ are the proper transforms of $C$ and $B$, then $(C^{\prime}\cdot B^{\prime})=(C\cdot B)-m_{B}m_{C}.$ Moreover, $(C\cdot B)_{q}=m_{B}m_{C}$ if and only if $C$ and $B$ are transverse at $q$, i.e. their tangent directions at $p$ are different.
2.2. Hypertangency at unibranched points of plane curves
Let $B$ now be an integral plane curve and $q$ a smooth point of $B$; we describe curves that are hypertangent to $B$ at $q$. See Subsection 4.1 for some examples.
Theorem 2.2.1.
Let $B,C\subset\mathbb{P}^{2}$ be two integral curves of degree at least $2$ such that $B\cap C=\{q\}$ and $q$ is a unibranched point for $B$ and $C$. If $q$ is a $(1,l)$-point for $B$, then $q$ is an $(m,lm)$-point of $C$ for some $m\geq 1$. Moreover the $\delta$ invariant of $q$ on $C$ satisfies
$\delta_{C}(q)\geq(m-1)(\deg B\deg C-m)/2.$
Proof.
Set $b=\deg B$, $d=\deg C$, and let $q$ be an $(m,n)$ point of $C$. Hence $l\leq b$ and $1\leq m<n\leq d$. By hypothesis,
$(C\cdot B)_{q}=(C\cdot B)=bd>m$
therefore $C$ and $B$ must have the same tangent line at $q$ (by Remark 2.1.2); we write $L$ for this line. Let $C^{1}$, $B^{1}$ and $L^{1}$ be the proper transforms of $C,B$ and $L$ in the blow-up, $S^{1}$, of $\mathbb{P}^{2}$ at $q$, and let $q^{1}\in C^{1}$ be the point lying over $q$. By hypothesis $C^{1}$ and $B^{1}$ meet only in $q^{1}$, hence (as $d>m$, $b\geq 2$)
$(C^{1}\cdot B^{1})_{q^{1}}=(C^{1}\cdot B^{1})=bd-m>bm-m\geq 2m-m=m.$
Now, $B^{1}$ is smooth at $q^{1}$ and ${\operatorname{mult}}_{q^{1}}(C^{1})\leq{\operatorname{mult}}_{q}(C)=m$, hence
$(C^{1}\cdot B^{1})_{q^{1}}>m\geq{\operatorname{mult}}_{q^{1}}(C^{1}){\operatorname{mult}}_{q^{1}}(B^{1}),$
hence $C^{1}$ and $B^{1}$ are tangent in $q^{1}$.
We now prove that $C$ has an $(m,lm)$-point at $q$. Suppose $l=2$, so that $q$ is a $(1,2)$-point of $B$. Lemma 2.1.1 implies that $B^{1}$ is neither tangent to $L^{1}$, nor to the exceptional divisor, hence the same holds for $C^{1}$. Hence, by the same Lemma, $q$ is a $(m,2m)$ point for $C$, and we are done.
To set-up an inductive argument we introduce a sequence of blow-ups as follows. We already considered the blow-up, $S^{1}$, of $S^{0}=\mathbb{P}^{2}$ at $q^{0}=q$. For $i\geq 1$ we let $S^{i}$ be the blow-up of $S^{i-1}$ at the unique point $q^{i-1}\in C^{i-1}$ lying over $q$. We denote by $C^{i}$, $B^{i}$ and $L^{i}$ the proper transforms of $C,B$ and $L$. Notice that the multiplicity of $C^{i}$ at $q^{i}$ is at most $m$.
Claim.
Suppose $l\geq 3$; for every $21\leq i\leq l-$ 2 the curve $C^{i}$ has an $(m,n-im)$-point at $q^{i}$, and $C^{i}$ and $B^{i}$ are tangent to $L^{i}$ in $q^{i}$.
We prove the claim by induction. The base is $i=1$. Now $q$ is a $(1,l\geq 3)$-point for $B$, hence Lemma 2.1.1 gives that $B^{1}$ is tangent to $L^{1}$ over $q$. We proved earlier that $B^{1}$ and $C^{1}$ are tangent in $q^{1}$, hence $C^{1}$ is tangent to $L^{1}$ at $q^{1}$; hence, by Lemma 2.1.1, $q$ is an $(m,n)$-point for $C$ with $n>2m$ and $C^{1}$ has a $(m,n-m)$-point. The proof of the base is complete.
To continue with the induction, suppose $C^{i-1}$ has a $(m,m-(i-1)n)$-point at $q^{i-1}$, and $B^{i-1}$ and $C^{i-1}$ are tangent to $L^{i-1}$ at $q^{i-1}$. Now $B^{i-1}$ has a $(1,l-(i-1))$-point in $q^{i-1}$; as $i\leq l-2$ we have $l-(i-1)\geq l-l+3=3$, hence $B^{i}$ is tangent to $L^{i}$. Now, $l\leq b$, hence $i\leq l-2\leq b-2$. Therefore
$(C^{i}\cdot B^{i})_{q^{i}}=(C^{i}\cdot B^{i})=bd-im>bm-im\geq bm-(b-2)m=2m>m.$
Hence $C^{i}$ and $B^{i}$ are tangent in $q^{i}$ at $L^{i}$. Therefore $C^{i}$ is tangent to $L^{i}$ and case (b) of Lemma 2.1.1 must occur for $C^{i-1}$, i.e. $m-(i-1)n>2m$ and $C^{i}$ has a $(m,n-im)$-point. The proof of the claim is complete.
Thus $C^{l-2}$ has an $(m,n-(l-2)m)$-point at $q^{l-2}$, and $B^{l-2}$ and $C^{l-2}$ are tangent to $L^{l-2}$ at $q^{l-2}$. Now, $B^{l-2}$ has a $(1,2)$-point at $q^{l-2}$, hence its proper transform, $B^{l-1}$, is neither tangent to $L^{l-1}$ nor to the exceptional divisor. We have, as $m<d$ and $l\leq b$
$(C^{l-1}\cdot B^{l-1})_{q^{l-1}}=(C^{l-1}\cdot B^{l-1})=bd-(l-1)m>bm-(b-1)m=m,$
hence $C^{l-1}$ and $B^{l-1}$ are tangent in $q^{l-1}$, hence $C^{l-1}$ is neither tangent to $L^{l-1}$ nor to the exceptional divisor. Now case (c) of Lemma 2.1.1 occurs for $C^{l-2}$, i.e. $q^{l-2}$ is a $(m,2m)$-point. Therefore $2m=n-(l-2)m$, hence $n=lm$, as stated.
Let us now study the $\delta$-invariant for $q$ as a point of $C$. Set
$h=\lceil(bd-m)/m\rceil.$
We prove, by induction on $i$, that $q^{i}$ is an $m$-fold point of $C^{i}$ for every $i\leq h-1$, and $B^{i}$ and $C^{i}$ are tangent in $q^{i}$. We already proved this for every $i\leq l-1$, hence the base case is settled and we assume $i\geq l$. The argument is similar to the one used in the previous part. Assume ${\operatorname{mult}}_{q^{i-1}}(C^{i-1})=m$, and $B^{i-1}$ tangent to $C^{i-1}$ in $q^{i-1}$. Then, as $i\leq h-1$, we have
$(C^{i}\cdot B^{i})_{q^{i}}=(C^{i}\cdot B^{i})=bd-im\geq bd-(h-1)m=$
$bd-(\lceil(bd-m)/m\rceil-1)m>bd-(bd/m-1)m=bd-bd+m=m$
(as $\lceil(bd-m)/m\rceil<bd/m$). Hence $(C^{i}\cdot B^{i})_{q^{i}}>m$, hence $C^{i}$ and $B^{i}$ are tangent at $q^{i}$. Now, the curve $B^{i-1}$ has a smooth point at $q^{i-1}$, hence $B^{i}$ is not tangent to the exceptional divisor in $q^{i}$. Hence the same holds for $C^{i}$. Hence case (a) of Lemma 2.1.1 does not occur for $C^{i-1}$, hence $q^{i}$ is an $m$-fold point of $C^{i}$. So we are done.
Since $C^{i}$ has an $m$-fold point at $q^{i}$ for every $i=0,\ldots,h-1$, by (2) the $\delta$-invariant of $q\in C$ satisfies
$\delta_{C}(q)\geq hm(m-1)/2=\lceil(bd-m)/m\rceil m(m-1)/2$
$\geq((bd-m)/m)m(m-1)/2=(bd-m)(m-1)/2.$
The proof is now complete. ∎
3. Hyper-bitangency for 3C-curves
In this section we concentrate on the case $S=\mathbb{P}^{2}$.
3.1. Definition and simple cases.
We study hyper-bitangent curves to a curve $B\subset\mathbb{P}^{2}$ which is the transverse union of three integral curves.
Definition 3.1.1.
A 3C-curve is a reduced plane curve $B=B_{1}\cup B_{2}\cup B_{3}$, with $B_{i}$ integral of degree $b_{i}\geq 1$, such that every point in $B_{i}\cap B_{j}$ is a node of $B$ for all $i\neq j$. We always assume $b_{1}\leq b_{2}\leq b_{3}$. We set
$B_{i}\cap B_{j}=\{p_{i,j}^{t},\quad t=1,\ldots,b_{i}b_{j}\}$
with $p^{t}_{i,j}=p^{t}_{j,i}$; we often omit the superscript $t$. Notice that the components of $B$ meet only pairwise, and transversally. We write $N:=\cup_{i\neq j}B_{i}\cap B_{j}.$
We begin with the case $b_{1}=b_{2}=b_{3}=1$. This is a particularly simple curve which can be easily handled.
Proposition 3.1.2.
Let $B=B_{1}\cup B_{2}\cup B_{3}$ be a 3C-curve of degree $3$. Then $\dim\operatorname{Hyp}_{d}(B,2)\geq 1$ for every $d\geq 1$.
Proof.
The curve $B$ is the union of three lines; set $B_{i}\cap B_{j}=\{p_{i,j}\}$ so that $N=\{p_{1,2},p_{1,3},p_{2,3},\}$.
Suppose $d=1$; then the one-dimensional space of lines through $\{p_{i,j}\}$, with $B_{i}$ and $B_{j}$ removed, lies in $\operatorname{Hyp}_{1}(B,2)$, and these are the only elements of $\operatorname{Hyp}_{1}(B,2)$. Hence $\dim\operatorname{Hyp}_{1}(B,2)=1$.
Let $d\geq 2$ and $C\in\operatorname{Hyp}_{d}(B,2)$. Obviously $C\cap N\neq\emptyset$, say $p_{1,2}\in C$. Then $C$ must be transverse to at least one between $B_{1}$ and $B_{2}$, say $C$ transverse to $B_{1}$, hence $C$ meets $B_{1}$ in a further point. Since $C$ must also meet $B_{3}$ we get $p_{1,3}\in C$, and $C$ cannot be transverse to $B_{3}$. We derive that $C$ is hypertangent to $B_{2}$ and $B_{3}$ respectively at $p_{1,2}$ and $p_{1,3}$. Now, setting $m=m_{p_{1,2}}(C)$ and $n=m_{p_{1,3}}(C)$, we have
$d=(C\cdot B_{1})=(C\cdot B_{1})_{p_{1,2}}+(C\cdot B_{1})_{p_{1,3}}=m+n\leq d$
hence $d=n+m$. Interchanging the three components, we derive
$\operatorname{Hyp}_{d}(B,2)=\bigcup_{m=1}^{d-1}\bigcup_{\stackrel{{\scriptstyle\quad i,j,h=1,2,3}}{{i\neq j}}}\operatorname{Hyp}_{d}^{m}(B_{i};p_{h,i})\cap\operatorname{Hyp}_{d}^{d-m}(B_{j};p_{h,j}).$
Consider $m=1$ and the subspace $\operatorname{Hyp}_{d}^{1}(B_{i};p_{h,i})\cap\operatorname{Hyp}_{d}^{d-1}(B_{j};p_{h,j}).$ Now, $\operatorname{Hyp}_{d}^{1}(B_{i};p_{h,i})$ is the space of degree-$d$ curves passing through $p_{h,i}$, which is a linear subspace of codimension $d$ in $\mathbb{P}^{d(d+3)/2}$. And $\operatorname{Hyp}_{d}^{d-1}(B_{j};p_{h,j})$ is the space of degree-$d$ curves having a $(d-1,d)$-point at $p_{h,j}$ with tangent line equal to $B_{j}$. To compute its dimension we can assume that $p_{h,j}$ is the origin and the line $B_{j}$ has equation $y=0$. Then a curve $C\in\operatorname{Hyp}_{d}^{d-1}(B_{j};p_{h,j})$ has equation $\sum_{d-1\leq i+j\leq d}a_{i,j}x^{i}y^{j}=0$ with $a_{i,j}=0$ for every $i+j=d-1$ and $i\neq 0$. One easily checks that such polynomials form a subspace of codimension equal to $d(d-1)/2+d-1=(d^{2}+d-2)/2$ hence
$\operatorname{codim}\operatorname{Hyp}_{d}^{1}(B_{i};p_{h,i})\cap\operatorname{Hyp}_{d}^{d-1}(B_{j};p_{h,j})\leq d+(d^{2}+d-2)/2=(d^{2}+3d-2)/2.$
As $\dim\operatorname{Hyp}_{d}(B,2)\supset\operatorname{Hyp}_{d}^{1}(B_{i};p_{h,i})\cap\operatorname{Hyp}_{d}^{d-1}(B_{j};p_{h,j})$ we get
$\dim\operatorname{Hyp}_{d}(B,2)\geq d(d+3)/2-(d^{2}+3d-2)/2=1.$
3.2. Hyper-bitangent lines
We now describe hyper-bitangent lines to a 3C-curve of degree at least $4$. This is quite elementary, possibly part of it already known. We include it for completeness and lack of references.
Proposition 3.2.1.
Let $B=B_{1}\cup B_{2}\cup B_{3}$ be a 3C-curve of degree $b\geq 4$.
  1. (a)
    If $b=4$ then $|\operatorname{Hyp}_{1}(B,2)|=6$.
    More precisely, $C\in\operatorname{Hyp}_{1}(B,2)$ if and only if $C$ is one of the four lines through $B_{3}\cap(B_{1}\cup B_{2})$ different from $B_{1}$ and $B_{2}$, or $C$ is one of the two lines tangent to $B_{3}$ and passing through the point $B_{1}\cap B_{2}$.
[…]
  2. (b)
    If $b\geq 5$ then $\operatorname{Hyp}_{1}(B,2)$ is finite, and it is empty if $B$ is general.
    More precisely, if $b_{1}=b_{2}=1$ then $|\operatorname{Hyp}_{1}(B,2)|\leq 3b_{3}(b_{3}-2)$, and $|\operatorname{Hyp}_{1}(B,2)|\leq 2|N|$ otherwise.
Proof.
If $b=4$ then $b_{1}=b_{2}=1$ and $b_{3}=2$; set $B_{1}\cap B_{2}=\{p_{1,2}\}$. We first look at the lines through two points of $N$. If $C$ is a line through $p_{1,3}^{t}$ and $p_{2,3}^{t^{\prime}}$ it clearly lies in $\operatorname{Hyp}_{1}(B,2)$. If $C$ is a line through $p_{i,3}^{t}$ and $p_{i,3}^{t^{\prime}}$ then it is equal to $B_{i}$ which is not possible. Now suppose that $C\in\operatorname{Hyp}_{1}(B,2)$ is not one of these lines; then we must have $p_{1,2}\in C$ and $C$ must meet $B_{3}$ in a unique point, hence $C$ must be tangent to $B_{3}$. Part (a) is proved.
Let $b\geq 5$ and let $B$ a general curve; we can make the following assumptions. If $b_{3}\geq 3$ then $B_{3}$ has finitely many flexes, hence finitely many flex lines, and finitely many bitangent lines; we assume that $B_{1}$ and $B_{2}$ do not pass through any flex of $B_{3}$ or through any point where $B_{3}$ meets its bitangent lines, and that $B_{1}\cap B_{2}$ intersects no line hyper-bitangent to $B_{3}$.
If $b_{3}=2$ then $b_{2}=2$, we assume that $B_{1}$ does not intersect $B_{2}\cup B_{3}$ at any point where $B_{2}\cup B_{3}$ meets its bitangent lines.
Finally, given any $p_{2,3}\in B_{2}\cap B_{3}$, there are finitely many lines through $p_{2,3}$ that are tangent to $B_{3}$, hence there are finitely many points $r\in B_{3}$ such that the tangent line to $B_{3}$ at $r$ passes through $B_{2}\cap B_{3}$; we assume that $B_{1}$ does not intersect $B_{3}$ in any of these points. Similarly (if $b_{2}\neq 1$) we assume that $B_{1}$ does not intersect $B_{2}$ in any point lying on some tangent line to $B_{2}$ passing through a point in $B_{2}\cap B_{3}$.
By contradiction, let $C\in\operatorname{Hyp}_{1}(B,2)$. Suppose $|C\cap N|=2$.
If $b_{3}\geq 3$ then, as $N$ contains no flex of $B_{3}$, we have $C\cap B=\{p_{1,3},p_{2,3}\}$ and, as $C$ cannot be a bitangent to $B_{3}$, we have
$b_{3}=(C\cdot B_{3})_{p_{1,3}}+(C\cdot B_{3})_{p_{2,3}}\leq 3$
hence $b_{3}=3$. Now $C$ must be tangent to $B_{3}$ in one of the points, $p_{i,3}$. Hence $C$ is a line through $B_{2}\cap B_{3}$ tangent to $B_{3}$ and intersecting $B_{1}\cap B_{3}$; we excluded the existence of such lines, so we are done.
Let $b_{3}=2$, hence $b_{2}=2$. If $b_{1}=1$, up to switching $B_{2}$ and $B_{3}$ we have only the case $C\cap B=\{p_{1,2},p_{2,3}\}$. Clearly $C$ is tangent to $B_{3}$ at $p_{2,3}$ and to $B_{2}$ at $p_{1,2}$, hence it is a bitangent of $B_{2}\cup B_{3}$, intersecting $B_{1}\cap B_{2}$, which we excluded. Let $b_{1}=2$. Our generality assumptions prevent us from switching $B_{1}$ with $B_{2}$ or $B_{3}$, so we have more cases. If $C\cap B=\{p_{1,2},p_{1,3}\}$ then $C$ is a bitangent of $B_{2}\cup B_{3}$, which is excluded. If $C\cap B=\{p_{1,2},p_{2,3}\}$ then $C$ is tangent to $B_{3}$ at $p_{2,3}$ and intersects $B_{1}\cap B_{2}$, which is excluded. As we can switch $B_{2}$ with $B_{3}$ we are done. We thus proved that $|C\cap N|=1$.
Suppose $C\cap N=\{p_{1,2}\}$. Then $C$ meets $B_{3}$ in a point $r\not\in N$, and it is hypertangent to $B_{3}$ at $r$. If $b_{3}\geq 3$ then $r$ is a flex and, by our assumptions, the flex line does not pass through $B_{1}\cap B_{2}$. If $b_{3}=2$ then $b_{2}=2$ and $C$ is a bitangent of $B_{2}\cup B_{3}$, which is also excluded.
Suppose $C\cap N=\{p_{1,3}\}$, then either $b_{3}\geq 3$ and $C$ is a flex line of $B_{3}$ which is excluded, or $b_{3}=b_{3}=2$ and $C$ is a bitangent of $B_{2}\cup B_{3}$, which is excluded.
Suppose $C\cap N=\{p_{2,3}\}$. Now, $C$ cannot be tangent to both $B_{2}$ and $B_{3}$, hence $b_{2}=1$, hence $b_{1}=1$, hence $b_{3}\geq 3$, hence $p_{2,3}$ is a flex of $B_{3}$. A contradiction. We thus proved that $\operatorname{Hyp}_{1}(B,2)$ is empty for $B$ general.
If $B$ is an arbitrary curve, the proof shows that for a curve $C\in\operatorname{Hyp}_{1}(B,2)$ only two cases can occur. First case: $C$ is tangent to some component of $B$ in a point of $N$; since at each point of $N$ there are such tangent lines we have at most $2|N|$ possibilities for such a $C$.
Second case: $C$ meets all components of $B$ transversally along $N$. Then one easily checks that $b_{1}=b_{2}=1$, moreover $C$ passes through $B_{1}\cap B_{2}$ and is hypertangent to $B_{3}$ in a flex (or rather a hyper-flex) or in a unibranched singular point. It is well known that the number of flexes of $B_{3}$ is at most equal to $3b_{3}(b_{3}-2)$; hence in the second case we have at most $3b_{3}(b_{3}-2)$ possibilities for $C$.
If $b_{1}=b_{2}=1$ the first case occurs only with $C$ hypertangent $B_{3}$, hence the bound $\operatorname{Hyp}_{1}(B,2)\leq 3b_{3}(b_{3}-2)$ holds. ∎
3.3. Hyper-bitangent curves of higher degree
We begin with a geometric description of hyper-bitangent curves.
Theorem 3.3.1.
Let $d\geq 2$. Let $B=B_{1}\cup B_{2}\cup B_{3}$ be a 3C-curve of degree $b\geq 4$ such that $\operatorname{Hyp}_{d}(B,2)$ is not empty.
Then $b_{1}=1$$b_{2}\leq 2$ and the following occur.
  1. (1)
    If $b_{2}=1$, setting $B_{1}\cap B_{2}=\{p\}$, we have
    $\operatorname{Hyp}_{d}(B,2)=\bigcup_{\stackrel{{\scriptstyle i=1,2}}{{q\in B_{i}\cap B_{3}}}}\operatorname{Hyp}_{d}^{d-1}(B_{i};p)\cap\operatorname{Hyp}_{d}(B_{3};q).$
[…]
  2. (2)
    If $b_{2}=2$ then $\operatorname{Hyp}_{d}(B,2)$ is empty for $d\geq 3$ and
    $\operatorname{Hyp}_{2}(B,2)=\bigcup_{\stackrel{{\scriptstyle p\in B_{1}\cap B_{2}}}{{q\in B_{1}\cap B_{3}}}}\operatorname{Hyp}_{2}(B_{2},p)\cap\operatorname{Hyp}_{2}(B_{3},q).$
  3. (3)
    Every $C\in\operatorname{Hyp}_{d}(B,2)$ is rational (hence $\mathcal{E}(B)=\operatorname{Hyp}(B,2)$, with $\mathcal{E}(B)$ defined in (1)).
Proof.
Since $b\geq 4$ we have $b_{3}\geq 2$. Let $C\in\operatorname{Hyp}_{d}(B,2)$, then $|C\cap B|\leq 2$; on the other hand $B_{1}\cap B_{2}\cap B_{3}=\emptyset$, hence $|C\cap B|=2$.
Let us prove that $C\cap B_{3}\subset N$ and $|C\cap B_{3}|=1$. By contradiction, suppose $C\cap B_{3}$ contains a point not in $N$. Hence $C$ must intersect $B_{1}\cup B_{2}$ in exactly one point, $p_{1,2}\in B_{1}\cap B_{2}$. Now, $p_{1,2}$ is a unibranched $n$-fold point of $C$, for some $n<d$ (as $d>1$). Since $B_{1}$ and $B_{2}$ meet transversally, $C$ is tangent to one of them and transverse to the other, say $C$ is transverse to $B_{i}$ with $i<3$. Therefore
$b_{i}d=(C\cdot B_{i})=(C\cdot B_{i})_{p_{1,2}}=n<d$
a contradiction. Hence $C\cap B_{3}\subset N$.
By contradiction, suppose $|C\cap B_{3}|=2$; as $C$ must intersect $B_{1}$ and $B_{2}$ we have $C\cap B_{3}=\{p_{1,3},p_{2,3}\}\subset N$, with $p_{i,3}\in B_{i}\cap B_{3}$. Since $B_{3}$ and $B_{i}$ meet transversally, $C$ must be transverse to either $B_{i}$ or $B_{3}$ at $p_{i,3}$. If $C$ is transverse to $B_{i}$ then, arguing as above, we get $(C\cdot B_{i})<d$, a contradiction. Hence $C$ is transverse to $B_{3}$ at both $p_{1,3}$ and $p_{2,3}$. Set $m_{i}={\operatorname{mult}}_{p_{i,3}}(C)$, so that $m_{1}+m_{2}\leq d$. Then
$b_{3}d=(C\cdot B_{3})=(C\cdot B_{3})_{p_{1,3}}+(C\cdot B_{3})_{p_{2,3}}=m_{1}+m_{2}\leq d$
which is impossible, as $b_{3}\geq 2$.
We thus proved that $C\cap B_{3}=\{p_{i,3}\}$ for one $i\in\{1,2\}$. Hence $C$ is hypertangent to $B_{3}$ at $p_{i,3}$. As $C$ is not transverse to $B_{3}$ a $p_{i,3}$, it must be transverse to $B_{i}$, hence it must meet $B_{i}$ in a further point. As $C$ must meet the other component, $B_{j}$ with $j\neq i,3$, there exists a point $p_{1,2}\in B_{1}\cap B_{2}$ such that $p_{1,2}\in C$. We obtain $C\cap B=\{p_{1,2},p_{i,3}\}$ and $C\cap B_{j}=\{p_{1,2}\}$. Now $B_{j}$ and $C$ meet only at $p_{1,2}$, hence $C$ cannot be transverse to $B_{j}$ at this point, hence it must be transverse to $B_{i}$. Therefore
$b_{i}d=(C\cdot B_{i})=(C\cdot B_{i})_{p_{1,2}}+(C\cdot B_{i})_{p_{i,3}}={\operatorname{mult}}_{p_{1,2}}(C)+{\operatorname{mult}}_{p_{i,3}}(C)\leq d$
hence $b_{i}=1$ and equality must hold, i.e. ${\operatorname{mult}}_{p_{1,2}}(C)+{\operatorname{mult}}_{p_{i,3}}(C)=d.$ Therefore $b_{1}=1$ and $C\in\operatorname{Hyp}_{d}^{d-m}(B_{2},p_{1,2})\cap\operatorname{Hyp}_{d}^{m}(B_{3},p_{1,3})$, with $m={\operatorname{mult}}_{p_{1,3}}(C)$.
We now show that $m=1$. Since $B_{3}$ has degree $\geq 2$ and is smooth at $p_{1,3}$, we can apply Theorem 2.2.1, which implies that $p_{1,3}$ is an $(m,lm)$ point for some $l\geq 2$. Suppose $m\geq 2$; by the same theorem, the $\delta$-invariant of $C$ at $p_{1,3}$ is $\geq(b_{3}d-m)(m-1)/2$. On the other hand $C$ has also a $(d-m)$-fold point at $p_{1,2}$, hence the $\delta$ invariant at $p_{1,2}$ is $\geq(d-m)(d-m-1)/2$. Therefore the genus of $C^{\nu}$ satisfies
$g(C^{\nu})\leq{\binom{d-1}{2}}-(b_{3}d-m)(m-1)/2-(d-m)(d-m-1)/2=$
${\binom{d-1}{2}}-(d^{2}-d-2md+b_{3}md-b_{3}d+2m)/2=$
${\binom{d-1}{2}}-(d^{2}+d(-1+m(b_{3}-2)-b_{3})+2m)/2\leq$
${\binom{d-1}{2}}-(d^{2}-3d+4)/2=(d^{2}-3d+2)/2-(d^{2}-3d+4)/2<0$
as $m\geq 2$ and $b_{3}\geq 2$. This is impossible. Hence $m=1$ and
(4) $C\in\operatorname{Hyp}_{d}^{d-1}(B_{2},p_{1,2})\cap\operatorname{Hyp}_{d}(B_{3},p_{1,3}).$
If $b_{2}=1$ we can switch roles between $B_{1}$ and $B_{2}$; part (1) is proved.
Assume $b_{2}\geq 2$. Then, as $B_{2}$ is smooth at $p_{1,2}$, we can apply Theorem 2.2.1, which gives that $C$ has a $(n,ln)$ point at $p_{1,2}$, for some $n\geq 1$. Now (4) implies that $p_{1,2}$ is a $(d-1,d)$-point of $C$. Therefore $(n,ln)=(d-1,d)$, hence $n=1$ and $d=2$. Therefore $\operatorname{Hyp}_{d}(B,2)$ is empty if $d\geq 3$, and the proof of part (2) is complete.
A curve of degree $d$ having a $(d-1)$-fold point is necessarily rational, hence part (3) follows from the previous parts. ∎
If $d\geq 2$, Theorem 3.3.1 implies that $\operatorname{Hyp}_{d}(B,2)=\emptyset$ whenever $b_{1}>1$, or $b_{2}>3$, or $b_{2}=2$ and $d\geq 3$. We now treat the remaining cases. We denote by $\operatorname{Hyp}_{\geq 2}(B,2)$ the set of curves of degree at least $2$ that are hyper-bitangent to $B$, i.e. $\operatorname{Hyp}_{\geq 2}(B,2):=\cup_{d\geq 2}\operatorname{Hyp}_{d}(B,2)$.
Theorem 3.3.2.
Let $B$ be a 3C-curve such that $(b_{1}=1,b_{2}\leq 2,b_{3}\geq 2)$.
  1. (a)
    If $(b_{1},b_{2},b_{3})=(1,1,b_{3})$ then $|\operatorname{Hyp}_{\geq 2}(B,2)|\leq 2b_{3}$.
  2. (b)
    If $(b_{1},b_{2},b_{3})=(1,2,b_{3})$ then $|\operatorname{Hyp}_{\geq 2}(B,2)|\leq b_{3}$.
  3. (c)
    In both cases above, if $B$ is general then $|\operatorname{Hyp}_{\geq 2}(B,2)|=0$.
Proof.
Let $C\in\operatorname{Hyp}_{d}(B,2)$; by Theorem 3.3.1, up to switching $B_{1}$ and $B_{2}$ when $b_{2}=1$, we have $C\in\operatorname{Hyp}_{d}^{d-1}(B_{2};p)\cap\operatorname{Hyp}_{d}(B_{3};q)$ with $p\in B_{1}\cap B_{2}$ and $q\in B_{1}\cap B_{3}.$ Hence $C$ has a $(d-1,d)$-fold point at $p$ where it is tangent to $B_{2}$, and a smooth point at $q$ where it is tangent to $B_{3}$. We can choose homogeneous coordinates $(X,Y,Z)$ in $\mathbb{P}^{2}$ so that $p=(0:1:0)$ and the tangent line to $C$ at $p$ has equation $z=0$. Therefore in the open subset where $Y\neq 0$ the curve $C$ has affine equation $z^{d-1}=\sum_{i=0}^{d}c_{i}x^{i}z^{d-i}.$ We can assume that $q=(0:0:1)$ and the tangent line to $B_{3}$ has equation $y=0$. Hence $c_{0}=c_{1}=0$ and the affine equation of $C$ where $Z\neq 0$ is
$y=g(x)\quad\quad\text{ where }\quad\quad g(x):=c_{d}x^{d}+c_{d-1}x^{d-1}+\ldots+c_{2}x^{2}.$
Claim.
$g(x)=c_{d}x^{d}$.
We assume $d\geq 3$ (otherwise it is obvious). Let $f(x,y)=0$ be the affine equation of $B_{3}$; as $B_{3}$ is smooth at $q$ and tangent to $y=0$, we have
(5) $f(x,y)=y+\sum_{2\leq i+j\leq n}a_{i,j}x^{i}y^{j}$
with $n=b_{3}$ for simplicity. We have $(C\cdot B_{3})=(C\cdot B_{3})_{q}=dn$, therefore $f(x,g(x))=\lambda x^{dn}$ for some $\lambda\neq 0$. Hence the following identity holds
(6) $g(x)+\sum_{2\leq i+j\leq n}a_{i,j}x^{i}g(x)^{j}=\lambda x^{dn}.$
Now, for every $i,j$ as above, the product $x^{i}g(x)^{j}$ is a sum of monomials in $x$ whose degrees range in the set $I(i,j)$, where
$I(i,j)=[i+2j,i+jd]\cap\mathbb{N}.$
We have
$i+dj\leq d(i+j)\leq dn$
with equality if and only if $j=n$ and $i=0$. We have
$I(0,n)=[2n,dn],\quad\quad I(1,n-1)=[2n-1,dn-d+1],$
therefore in the left side of (6), the following monomials (up to scalar)
$x^{dn},\ x^{dn-1},\ldots,x^{dn-d+2}$
appear only in $g(x)^{n}$. For (6) to hold, in its left side
(a) the coefficient of $x^{dn}$ is non zero, hence $a_{0,n}\neq 0$ and $c_{d}\neq 0$;
(b) the coefficients of $x^{dn-1},\ldots,x^{dn-(d-2)}$ are zero.
We now show, by induction on $h$, that (a) and (b) imply that $c_{d-h}=0$ for $h=1,\ldots,d-2$. We write $g(x)^{n}$ as follows
$g(x)^{n}=\sum_{k=1}^{d-2}\Bigg{(}\sum_{\stackrel{{\scriptstyle 2\leq i_{1}<\ldots<i_{k}\leq d}}{{\sum_{1}^{k}n_{j}=n,\ n_{j}\geq 1}}}\mu_{n_{1},\ldots,n_{k}}c_{i_{1}}^{n_{1}}\cdot\ldots\cdot c_{i_{k}}^{n_{k}}x^{n_{1}i_{1}+\ldots+n_{k}i_{k}}\Bigg{)}$
where $\mu_{n_{1},\ldots,n_{k}}$ are positive integers which we can ignore. Since $i_{k}\leq d$ and $i_{j}<i_{j+1}$ we have $i_{j}\leq i_{k}-(k-j)\leq d-(k-j)$, hence the exponent of $x$ above satisfies the following
$\sum_{j=1}^{k}n_{j}i_{j}$ $\leq\sum_{j=1}^{k}n_{j}(d-(k-j))=n_{1}(d-(k-1))+\ldots+n_{k-1}(d-1)+n_{k}d$
$=d\sum_{j=1}^{k}n_{j}-\sum_{j=1}^{k-1}n_{j}(k-j)\leq dn-(k-1)k/2$
where in the second inequality we used $n_{j}\geq 1$ for all $j\leq k-1$. Hence we have equality if and only if $i_{j}=d-(k-j)$ for $j\leq k$ and $n_{j}=1$ for $j\leq k-1$. If these conditions are satisfied, we furthermore have
$\sum_{j=1}^{k}n_{j}i_{j}=dn-1\quad\text{ if and only if }\quad\ k=2,$
i.e. $i_{1}=d-1$, $i_{2}=d$, $n_{1}=1$, $n_{2}=n-1$. Therefore the term $x^{dn-1}$ appears in $g(x)^{n}$ only once, with coefficient equal to (a positive integer multiple of) $c_{d}^{n-1}c_{d-1}$. Hence $x^{nd-1}$ appears in the left of (6) with coefficient $a_{0,n}c_{d-1}c_{d}^{n-1}$; as this coefficient must be zero and $a_{0,n}c_{d}\neq 0$, we get $c_{d-1}=0$. The induction base is proved.
Assume $c_{d-1}=c_{d-2}=\ldots=c_{d-h+1}=0$. We have a non-zero coefficient of $x^{\sum_{1}^{k}n_{j}i_{j}}$ only if $i_{j}\not\in\{i_{d-1},i_{d-2},\ldots,i_{d-h+1}\}$, in which case we have
$\sum_{j=1}^{k}n_{j}i_{j}$ $\leq n_{1}(d-h-k+2)+\ldots+n_{k-2}(d-h-1)+n_{k-1}(d-h)+n_{k}d$
$=dn-h\sum_{j=1}^{k-1}n_{j}-\sum_{j=1}^{k-2}j\leq dn-h(k-1)-(k-1)(k-2)/2.$
The first inequality is an equality if and only if $i_{k}=d$, and $i_{k-1}=d-h$, and $i_{j}=i_{j+1}-1$ for $j<k-1$; the second inequality is an equality if and only if $n_{j}=1$ for all $j\leq k-1$. If these conditions hold, we furthermore have $\sum_{j=1}^{k}n_{j}i_{j}=dn-h$ if and only if $k=2.$ Therefore, arguing as above, $x^{dn-h}$ appears in $g(x)^{n}$ with coefficient $c_{d-h}c_{d}^{n-1}$. Hence it appears in the left of (6) with coefficient $a_{0,n}c_{d-h}c_{d}^{n-1}$; as this must be zero, we get $c_{d-h}=0$. The claim is proved.
Thus $C$ has equation $y=c_{d}x^{d}$, hence $q$ is a $(1,d)$-point of $C$. But $C$ is hypertangent to $B_{3}$ at $q$, hence Theorem 2.2.1 implies that $B_{3}$ also has a $(1,d)$-point at $q$.
Assume $d\geq 3$. If $B$ is general we can assume that no point in $B_{3}\cap(B_{1}\cup B_{2})$ is a a flex of $B_{3}$. Hence $q$ is a $(1,2)$-point of $B_{3}$ and we get a contradiction. Therefore $\operatorname{Hyp}_{d}(B,2)=\emptyset$ if $B$ is general.
If, instead, $B_{3}$ has a $(1,l)$-point in $q$ with $l\geq 3$, we get $d=l$, hence the degree of $C$ is determined by $q$. Moreover no term of type $x^{i}$ with $i<d$ can appear in the equation of $B_{3}$ (for $q$ is a $(1,d)$-point), hence $a_{i,0}=0$ for all $i<d$, and from (6) we derive
$c_{d}x^{d}+a_{d,0}x^{d}+(\text{terms of degree }>d)=\lambda x^{dn}.$
As $n\geq 2$ we get $c_{d}=-a_{d,0}$. This proves that $C$ (if it exists) is uniquely determined by $q$.
Assume $d=2$, then $q$ is a $(1,2)$-point for $C$ and $B_{3}$. Now (6) gives
$c_{2}x^{2}+a_{2,0}x^{2}+a_{1,1}c_{2}x^{3}+a_{3,0}x^{3}+(\text{terms of degree }>3)=\lambda x^{2n}.$
We obtain $a_{2,0}=-c_{2}$ and $a_{1,1}c_{2}=-a_{3,0}$, hence $a_{3,0}=a_{1,1}a_{2,0}$. Hence $B_{3}$ is not a general curve of degree $n$ (for its equation, (5), must satisfy $a_{3,0}=a_{1,1}a_{2,0}$). Hence $\operatorname{Hyp}_{2}(B,2)=\emptyset$ if $B$ is general. (c) is proved.
If $B_{3}$ is not general, then, as before, $C$ is determined by the condition $c_{2}=-a_{2,0}$, hence it is determined by $q$.
Summarizing, for all $d\geq 3$ we proved that, for every $q\in B_{1}\cap B_{3}$ there exists at most one curve $C\in\operatorname{Hyp}_{d}(B_{2};p)\cap\operatorname{Hyp}_{d}(B_{3};q)$; as $q$ varies in $B_{1}\cap B_{3}$ we get at most $b_{3}$ curves in $\operatorname{Hyp}(B,2)$. If $b_{2}=1$, by Theorem 3.3.1, the same argument applies by taking $q\in B_{2}\cap B_{3}$ hence we might have $b_{3}$ new elements in $\operatorname{Hyp}(B,2)$. We thus proved that $|\operatorname{Hyp}_{\geq 2}(B,2)|\leq 2b_{3}$ if $b_{2}=1$ and $|\operatorname{Hyp}_{\geq 2}(B,2)|\leq b_{3}$ if $b_{2}=2$. ∎
For an example where the bound in the theorem is attained, see subsection 4.2. An immediate consequence is the following.
Theorem 3.3.3.
The set $\operatorname{Hyp}(B,2)$ of a 3C-curve $B$ of degree at least $5$ is finite, and it is empty if $B$ is general.
Proof.
If $d=1$ the statement follows from Proposition 3.2.1. If $d\geq 2$ and $b_{1}>1$ or $b_{2}>2$ it follows from Theorem 3.3.1. The remaining cases follow from Theorem 3.3.2.∎
4. Examples
We collect here some examples and special cases related to our results.
4.1. Examples of hypertangent curves
Let $B,C\subset\mathbb{P}^{2}$ be two integral curves as in Theorem 2.2.1, i.e. they have degree at least 2, $B\cap C=\{q\}$ with $q$ a unibranched point for $B$ and $C$. The following examples show that there are cases in which $q$ is a $(1,l)$-point for $B$ and $Q$ is a non-smooth point of $C$, i.e. is a $(m,lm)$ point with $m\geq 2$. In particular, in this setting, the type of the point $q$ for $B$ is different from its type as a point of $C$.
(1) Consider the following two curves:
$B:y=x^{2}\qquad C:(y-x^{2})^{3}+y^{7}.$
One easily checks that $C\in\operatorname{Hyp}(B;q)$ where $q$ is the origin, and that $B$ has a $(1,2)$ point while $C$ has a $(3,6)$ point, so that $m=3$ and $l=2$ in our notation. More generally, the following pair of curves,
$B:y=x^{2}\qquad C:(y-x^{2})^{c}+y^{2c+1},$
gives similar examples where $B$ has a $(1,2)$ point and $C$ has a $(c,2c)$ point for every $c\geq 3$.
(2) This second example shows that a similar situation can happen when $l$ is not 2. Consider the following two curves:
$B:y=x^{3}\qquad C:(y-x^{3})^{3}+y^{9}.$
In this case one checks that $q$ is a $(1,3)$-point for $B$ while it is a $(3,9)$-point for $C$; as before $C\in\operatorname{Hyp}(B;q)$.
We stress that both these examples do not yield examples of curves in $\operatorname{Hyp}(\tilde{B},2)$ for a 3C-curve $\tilde{B}$ containing $B$. In fact, the proof of Theorem 3.3.2 shows that the curves $C$ described above will have to meet the curve $\tilde{B}$ in more than two points.
4.2. Explicit examples of hyper-bitangent conics
We show that the bounds in Theorem 3.3.2 are sharp, in the sense that there are examples in which the set $\operatorname{Hyp}_{\geq 2}(B,2)$ consists of exactly $2b_{3}$ curves.
Figure 1. Example of 4 conics in $\operatorname{Hyp}_{2}(B,2)$
For an explicit example consider the following three components of a 3C-curve $B=B_{1}\cup B_{2}\cup B_{3}$ with $(b_{1},b_{2},b_{3})=(1,1,2)$:
$B_{1}:x=0,\qquad B_{2}:z=0,\qquad B_{3}:zy=x^{2}-y^{2}.$
Let $\{q_{1},q_{2},q_{3},q_{4}\}=(B_{1}\cap B_{3})\cup(B_{2}\cap B_{3})$, where
$q_{0}=(0:0:1),\quad q_{1}=(0:-1:1),\quad q_{2}=(1:1:0),\quad q_{3}=(1:-1:0).$
Then for $i=0,1,2,3$ the following conics $C_{i}$ satisfy $C_{i}\in\operatorname{Hyp}_{2}(B_{3};q_{i})$:
$C_{0}:zy=x^{2},$ $C_{1}:zy=-z^{2}-x^{2},$
$C_{2}:4xz+8xy-8x^{2}=z^{2},$ $C_{3}:4xz+8xy+8x^{2}=-z^{2}.$
Moreover, $C_{0},C_{1}\in\operatorname{Hyp}_{2}(B_{2};p)$ and $C_{2},C_{3}\in\operatorname{Hyp}_{2}(B_{1};p)$ where $p=(0:1:0)$. In other words $B_{2}$ is the tangent line to $C_{0}$ and $C_{1}$ at $p$, while $B_{1}$ is the tangent line to $C_{2}$ and $C_{3}$ at $p$. Therefore, for every $i=0,1,2,3$, $C_{i}\in\operatorname{Hyp}_{2}(B;p,q_{i})$ and hence $\lvert\operatorname{Hyp}_{2}(B,2)\rvert=4$ reaching the upper bound of (b) in the statement of Theorem 3.3.2. A picture of these conics in the affine patch $y=1$ can be seen in figure 1.
4.3. Curves with many components
If a curve $B$ has more than three components, it is not hard to prove that the only hyper-bitangent curves to $B$ are lines, and describe such lines effectively. We include the analysis of this case for completeness.
Proposition 4.3.1.
Let $B$ be a reduced curve of degree $b$ with $c\geq 4$ irreducible components, such that every point in the intersection of two components is a node of $B$.
  1. (a)
    If $c\geq 5$ then $\operatorname{Hyp}(B,2)=\emptyset$.
  2. (b)
    If $d=1$ then $\operatorname{Hyp}_{1}(B,2)$ is finite; moreover
    1. (i)
      if $b=4$ then $|\operatorname{Hyp}_{1}(B,2)|=3$;
    2. (ii)
      if $B$ does not contain two lines, then $\operatorname{Hyp}_{1}(B,2)=\emptyset$;
    3. (iii)
      if $b\geq 5$ and $B$ is general then $\operatorname{Hyp}_{1}(B,2)=\emptyset$.
  3. (c)
    If $d\geq 2$ then $\operatorname{Hyp}_{d}(B,2)=\emptyset$.
Proof.
Write $B=B_{1}\cup\ldots\cup B_{c}$ with $c\geq 4$; by hypothesis through every node of $B$ there pass at most two components. Suppose there exists a curve $C\in\operatorname{Hyp}(B,2)$. Then $C$ meets $B$ in at most two points, and must intersect all components of $B$. Hence $C\cap B=\{p,q\}$ and $p$, $q$ belong to exactly two components, say $p\in B_{1}\cap B_{2}$ and $q\in B_{3}\cap B_{4}$. In particular, $B$ has only 4 components, proving (a).
Let $d=1$ and $C\in\operatorname{Hyp}_{1}(B,2)$. As we said, $C\cap B=\{p,q\}$. Since $B$ has finitely many (intersection) nodes, $\operatorname{Hyp}_{1}(B,2)$ is finite.
If $b=4$ then $B$ is the union of 4 lines. It is clear that $\operatorname{Hyp}_{1}(B,2)$ is made of the three bitangent lines of $B$ not contained in $B$ (namely the three lines not in $B$ and joining a pair of nodes of $B$).
Let $b\geq 5$. We can assume that one component, $B_{4}$, has degree $\geq 2$. Now, as $B_{3}$ and $B_{4}$ meet transversally in $q$ and $C\cap B_{4}=\{q\}$, the line $C$ is necessarily hypertangent to $B_{4}$ at $q$; also, $B_{3}$ must have degree $1$, for it meets $C$ transversally in only one point. Now, if such a line $C$ exists it is unique and has to pass through the point $p$ as well. Arguing in the same way for $B_{1}$ and $B_{2}$ we have that at least one between $B_{1}$ and $B_{2}$ has degree $1$. Hence for $\operatorname{Hyp}_{1}(B,2)$ to be non-empty at least two components of $B$ have degree $1$.
Now, if the curve $B$ is general, we can assume that no such line exists, i.e. we can assume that for every point $q\in B_{i}\cap B_{j}$ the tangent lines to $B_{i}$ and $B_{j}$ in $q$ do not pass through any other intersection node of $B$. This proves that $\operatorname{Hyp}_{1}(B)$ is empty if $B$ is general. (b) is proved.
Let $d\geq 2$. By contradiction, let $C\in\operatorname{Hyp}_{d}(B,2)$. As before, we assume $C\cap B=\{p,q\}$ with $p\in B_{1}\cap B_{2}$ and $q\in B_{3}\cap B_{4}$. Now, $B_{1}$ and $B_{2}$ meet transversally, hence $C$ in $p$ must be transverse to at least one between $B_{1}$ and $B_{2}$; say $C$ is transverse to $B_{1}$. Hence $C$ must intersect $B_{1}$ in a further point, and this point must be $q$, which is impossible as $q$ cannot belong to three components of $B$. ∎
4.4. Hypertangency of rational curves
For an integral curve $B\subset\mathbb{P}^{2}$ of degree $b\geq 4$ having at most nodal singularities, Conjecture 1 predicts that there exist only finitely many rational curves hyper-bitangent to $B$, i.e. the set $\mathcal{E}(B)$ is finite. We will provide an example showing the necessity, in the conjecture, that $B$ have only nodal singularities.
For every integer $b\geq 3$ we denote by $Q_{b}\subset\mathbb{P}^{2}$ the curve given by the homogeneous equation
$z^{b-1}y=x^{b}.$
The curve $Q_{b}$ is smooth at $q_{0}=(0:0:1)$, with tangent line, $L_{0}$, of equation $y=0$. We have $Q_{b}\cap L_{0}=\{q_{0}\}$, so $L_{0}$ is hypertangent to $Q_{b}$. Next, $Q_{b}$ has an $(b-1)$-fold unibranched point at $q_{\infty}=(0:1:0)$ with tangent line $L_{\infty}$ of equation $z=0$. We have $Q_{b}\cap L_{\infty}=\{q_{\infty}\}$ so that $L_{\infty}$ is hypertangent and $q_{\infty}$ is an $(b-1,b)$-point.
[…]
One checks easily that $Q_{b}$ is integral, has no other singular point, and is a rational curve; it is, of course, not a nodal curve.
Proposition 4.4.1.
For every $d\geq b$ and $b\geq 4$ we have
(7) $\dim\operatorname{Hyp}_{d}(Q_{b},1)\geq 1$
and
(8) $\dim\mathcal{E}(B)\geq 1$
Proof.
Consider the curve $C_{t}$ of equation
$y-x^{b}+ty^{d}=0$
with $t\in\mathbb{C}$. It is clear that $C_{t}$ lies in $\operatorname{Hyp}_{d}(Q_{b};q_{0})$ for every $t\neq 0$ hence (7) follows. Notice that $C_{t}$ is smooth, hence not rational, for $t\neq 0$.
To prove (8) we will exhibit a one-dimensional family of curves in $\mathcal{E}(B)$. For every $t\neq 0,1$ let $R_{t}$ be the curve having equation
$z^{b-1}y=tx^{b}.$
It is easy to check that $R_{t}$ is integral, and its singular locus consists of $q_{\infty}$ which is a $(b-1,b)$-singular point, therefore $R_{t}$ is a rational curve. Moreover, we have
$R_{t}\cdot Q_{b}=bq_{0}+b(b-1)q_{\infty}$
hence $R_{t}\in\operatorname{Hyp}(Q_{b};q_{0},q_{\infty})$. As $t$ varies in $\mathbb{C}$ the curves $R_{t}$ form a family with a non-integral member, for $t=0$, hence the family is not constant. Therefore, the $R_{t}$ form a one-dimensional family of curves in $\mathcal{E}(B)$. ∎